Zachary Company incurs annual fixed costs of $111,000. Variable costs for Zachary’s product are $21.00 per unit, and the sales price is $35.00 per unit. Zachary desires to earn an annual profit of $57,000.


Required
Use the contribution margin ratio approach to determine the sales volume in dollars and units required to earn the desired profit.
Answer #1

Answers :

Desired Sales Dollars = $ 420,000

Desired Sales Units = 12,000 Units

Contribution Margin Per unit s= Selling Price- Variable cost = 35-21 =14

Contribution Margin ratio = contribution Margin / selling Price =14/35 = 40%

Desired Units = ( Fixed Cost + Desired Profit ) / Contribution MArgin

= (111,000+57,000)/ 14 = 12,000 Units (Answer )

Desired Dollars = 12,000*35 =420,000 (Answer)